zarkam21 (zarkam21):

The owner of a bike shop sells unicycles and bicycles and keeps inventory by counting seats and wheels. One day, she counts 21 seats and 30 wheels. How many of each type of cycle are there? Choose an equation in slope-intercept form that expresses the fact the number of bicycles plus the number of unicycles is 1. A. u=-b+21 B. -u=b+21 C. 21/6=u D. b+u=21 E. b=u+21

OpenStudy (anonymous):

The first thing that you need to realize is what slope-intercept form is. Remember that it is y=mx+b. I always say one variable on the left, while everything else is on the right. After you do that, you need to set up some equations. Seats Each cycle has 1 seat. You know there are 21 seats. So \[u + b = 21\] Manipulate that equation into something that looks like slope-intercept form. \[u + b = 21\] \[u + b - b = 21 - b\] \[u = 21 - b\] Now just rearrange the terms on the right side. You have an answer. If you needed the other equation, it would be set up like \[u + 2b = 30\]

zarkam21 (zarkam21):

so A =)

zarkam21 (zarkam21):

so A =)

zarkam21 (zarkam21):

so A =)